FEL Upgrade Project Weekly Brief
October 30 - November 3, 2000


The machine operated very well this week for users doing polymer ablation, carbon nanotube fabrication, and protein dynamics.


FEL Project management made presentations to the SURA/DOE review team on site this week for the biannual review of Jefferson Lab’s institutional management.  Dave Shirley, former Director of Lawrence Berkeley Lab, and John McTague, former chief technical officer for Ford were members of the review team giving specific scrutiny to the FEL program.  George Neil represented the FEL project and presented a paper at this week’s annual symposium of the Directed Energy Professional Society held in Albuquerque.

WBS 3 (Beam Physics)

A review of PARMELA simulations indicated that previous 135 pC solutions assumed sector injection line dipoles rather than the desired rectangular bends, which are operationally preferable due to the large (8:1 - 21:1) dynamic range of the Upgrade FEL driver.  Analysis of the problem with rectangular dipoles is proceeding, with a decision on dipole geometry anticipated prior to Revision 1.1 of the driver design.  Should a return to sector dipoles be indicated, it will be accommodated in the Rev 1.1 issue.

WBS 4 (Injector)

The 2nd nitrogen implanted electrode in the field emission test system is baked, aligned and ready to start data taking.


We visited the vendor working on the He vessel (PHPK, Inc., Westerville, OH) for a first article inspection.  The first article was approved including the important weld procedures.

WBS 6 (RF)

Continued debug and checkout of Zone 4 interlocks and misc. wiring.

Zone 3 had no progress due to FEL operations.

Received 1 incomplete proposal from PSC to upgrade the Injector power supplies.  Their pricing was 33 % higher than their original estimates.  We are contacting them to understand exactly what they are proposing to do.

The requisition for the 600 kV power supply upgrade made it to Purchasing this week with the proper signatures.

WBS 8  (I & C)

Of the Optical Beam Position Monitor (OBPM) subsystems: quad pyro detector/preamp, power supply/driver, remote control flip attenuator, 60 Hz chopper phase reference, 32 channel sample & hold, ADC and software progress was made this week on the chopper phase reference distribution and the sample & hold. The 32 channel sample & hold is a simple but newly designed VME board, of three that were made the first was damaged possibly by a power supply problem but the second one was successfully
commissioned. A great deal of effort was put into resolving mW level signals at <100 microseconds, but for the first draft I must relax the requirements.

None of the four candidates that were interviewed for the coordinator position were qualified, we are now placing an ad in the local papers.  The position demands an understanding of component level electronics (caps, resistors, opamps, and gates) along with fair proficiency in AutoCad for creating system prints.

The RPF for the shielded beam viewers generated only one bid from the five packages that were sent out. The price came considerably less than what was estimated.  Another bonus is that the bid is from a local machine shop and will be easier to track progress.  The RFP for the BPMs will hit the street today to a total of eight shops with ample time to bid (1 month).  These are also hoped to come it at competitive prices.

Some progress has also been made on vacuum controls; drawing are being updated and are being prepared for manufacturing. While considering the quantities required for the upgrade, "CEBAF" may also purchase some to bolster their spares inventory.

WBS 9 (Transport)

With the good solid model of the GW (window frame style) and the almost completely detailed (but obsolete) GX package I am finally able to get realistic cost estimates for the entire set of 22 dipoles.  The vendors can estimate costs of the remaining magnets with an outline sketch and a comparison to the GX, GW or the DU dipole from the IR Demo.  We worked on the letter and drawing sets to send to our prospective vendors.

As part of preparing the above package, we found an existence proof of the ability to morph the 3 inch gap Injector Dipoles from the existing 2 inch DU drawings.  The identical conductor can be used and will fit if the coil height contains 8 conductors rather than the DUís 5.

Robin Wines was able to find the problem with her input file of the magnetic model of the GX.  The several hour computer run is about to start.

QX (3.125" Quad) Magnets
Prototype Fabrication: The vendor took another cut on the aluminum mock-up of the core quadrant.  It is being measured.

Budgetary Quote: A new revision of the core drawing (Vaughan) was produced to declare tolerancing on the assembled core rather than the individual core quadrants.  This drawing will be sent out for quotes.

Measurement Probe: We ordered the two candidate probe materials and have started getting the shop to test their machining and stability properties.

Mark Wiseman of the Engineering Department assigned Robin Wines and Al Guerra to reprise their IR Demo roles as caretakers of the air & nitrogen and the cooling water (LCW) piping respectively.  We still need to appoint a keeper of the cable tray.

WBS 10 (Wiggler)

Optical Klystron System
The modified core clamp hardware was received so next week we should be able to prove out that the gap between the cores can be held to the ±0.001 inch level.

Drawings:  details of manifold bracket were signed off and design of the extra pieces for the wiggler lead connections started.

GS (Dispersion Section) steel and copper conductor has been shipped to the vender.

Further down the road is the design for mounting the dispersion dipole on the optical klystron support tube.

WBS 11 (Optics)

On the optical cavity we have defined mounting hardware for the mirror test stand.

We positioned optical diagnostics in the Optics Control Room in order to better utilize the pyro electric array on the spectrometer.


The drive laser green output had fallen and some work was done to fix it, but these did not work.  Fortunately the quantum efficiency of the gun photo cathode worked so that it does not effect operations.

As noted above, user operations have gone well this week.  On Wednesday, despite a 1 hr break in operations to accommodate the Institutional Management Review team tour, the FEL accommodated three different user experiments at three different wavelengths.

Privacy and Security Notice